OpenStudy (anonymous):

find the polar coordinates r>0 and r<0; (-6 sqrt 2, -6 sqrt2)

OpenStudy (experimentx):

for given (x,y) your polar coordinates are (r,\(\theta\)) r = \( \sqrt{x^2+y^2}\) \( \theta \) = \( \arctan(\frac{y}{x})\)

OpenStudy (anonymous):

so i got \[\sqrt{-12+\left( -12 \right)}\]

OpenStudy (anonymous):

one of the 12 has to be positive but im not sure how to make it positive

OpenStudy (experimentx):

\( (-6\sqrt{2})^2\)

OpenStudy (experimentx):

= 72

OpenStudy (anonymous):

oh 72+72=144 \[\sqrt{144}=\pm12\]

OpenStudy (experimentx):

+12

OpenStudy (anonymous):

the book as (12, 5pi/4) and (-12, pi/4); where did the pi/4 and 5pi/4 come from?

OpenStudy (experimentx):

angle, look at that theta.

OpenStudy (experimentx):

also try to draw coordinate first.

OpenStudy (anonymous):

Im a little confused how can i draw the coordinate if i only have 12 as an answer?

OpenStudy (experimentx):